OpenStudy (anonymous):

solve each system by substitution Please show how you got it and also how you found Y... Thanks :) 7x+2y=-6 -14x-4y=-2

OpenStudy (anonymous):

7x=-2y-6 (set the first equation in terms of x) -2(-2y-6)-4y=-2 (plug in the top equation to the bottom, we know that 7 goes into -14 -2 times) 4y+12-4y=-2 (simplify, a negative multiplied by a negative equals a positive) 12=-2 (add the y's) we have a problem here, 12 does not equal to -2, so the system has no solution. :)

OpenStudy (anonymous):

Thanks :)